OpenStudy (anonymous):

Consider the line that passes through the points (0, 4) and (-1, 3). Below are two different equations. Identify the true statement concerning both equations. Equation #1 Equation #2 y - 4 = 1(x - 0) y - 3 = 1(x - 1) Only equation #1 represents the line that passes through the two given points. Only equation #2 represents the line that passes through the two given points. Both equation #1 and equation #2 represent the line that passes through the two given points. Neither equation #1 nor equation #2 represents the line that passes through the two given

OpenStudy (jdoe0001):

plug the values to each and find out :) lemme do say the 1st one (0, 4) and (-1, 3) #1st (0, 4) case (4)-4 = 1( (0) - 0 ) 0 = 0 # the equation holds (-1, 3) case (3) - 4 = 1( (-1) - 0 ) -1 = -1 # the equation holds so, what about the 2nd equation? :)

OpenStudy (anonymous):

I don't understand

OpenStudy (jdoe0001):

(0, 4) and (-1, 3) are just ordered pairs for (x, y)  so to see if the lines pass through there, just plug in the provided ordered pairs the left side will equal the right side if it does 0 = 0 if the plugging in of the values gives you something like 5 = 3 well, 5 \(\bf \ne\) 3, so we know that, thus the equation doesn't hold, thus those coordinates aren't "good values" for that equation, which means, the line really doesn't pass through them

OpenStudy (anonymous):

Oh okay. Thank You for helping!!! :)

OpenStudy (jdoe0001):

yw